# CloudHeatMap: Heatmap-Based Monitoring for Large-Scale Cloud Systems

### v1

@@ IV Discussion
V Threats to Validity
The primary validity concerns, classified as per [33, 34], are as follows.
Construct Validity
The tool was iteratively developed with feedback from the IBM Ops team and tested in production for two months, ensuring alignment with evaluation goals and mitigating potential threats to construct validity.
Internal Validity
Rigorous testing and verification of the data were employed to minimize experimental errors.
External Validity
The monitoring tool is not limited to one software system. While not tested on other products, it is designed to be generic and applicable to any LCS.
VI Conclusion
We proposed CloudHeatMap, a heatmap-based tool to monitor LCS health, using the IBM Cloud Console as a software under study. The tool represents a significant advance in Cloud monitoring by providing actionable insights in real-time, allowing operators to make informed decisions about scaling, re-architecting, and cost-saving measures. Its unique approach to visualizing both HTTP and non-HTTP interactions makes it adaptable to various Cloud environments. We believe CloudHeatMap  will interest practitioners and academics and contribute to monitoring complex large-scale software systems.
Moving forward, we will focus on deriving the fourth Golden Rule of Monitoring — saturation — to provide a more complete picture of system health. Aggregating caller-callee pairs into call chains for graph-based root cause analysis will further enhance the tool’s diagnostic capabilities.

### v2

@@ 4. Discussion
5. Threats to Validity
The primary validity concerns, classified as per (Yin, 2009; Wohlin et al., 2012), are:
Construct Validity. The tool was iteratively developed with feedback from the IBM Ops team and tested in production for two months, ensuring alignment with evaluation goals and mitigating potential threats to construct validity.
Internal Validity. Rigorous testing and verification of the data were employed to minimize experimental errors.
External Validity. The monitoring tool is not limited to one software system (see Section 2.3.3). While not tested on other products, it is designed to be generic and applicable to any LCS.
6. Conclusion
We proposed CloudHeatMap, a heatmap-based tool to monitor LCS health, using the IBM Cloud Console as a software under study. The tool represents an advance in Cloud monitoring by providing actionable insights in real-time, allowing operators to make informed decisions about scaling, re-architecting, and cost-saving measures. Its approach to visualizing both HTTP and non-HTTP interactions makes it adaptable to various Cloud environments. We believe CloudHeatMap  will interest practitioners and academics and contribute to monitoring complex large-scale software systems.
Moving forward, we will extend the tool to the fourth Golden Rule of Monitoring — saturation — to provide a more complete picture of system health. Furthermore, aggregating caller-callee pairs into call chains for graph-based root cause analysis will further enhance the tool’s diagnostic capabilities.